# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v1

@@ 4.5 Latent variable estimation
For all of the simulation scenarios discussed in Section 4.1, we evaluate the estimation capabilities of the different fitted models specified in Section 4.2. We show model performance in terms of absolute bias $|\text{Bias}(x_{post},x_{true})|$ of the latent posterior estimates $x_{post}$ with respect to their true values $x_{true}$ as well as $\text{SD}(x_{post})$ as a measure of posterior sharpness. Since mean-field VI is known to underestimate the posterior SD (Kucukelbir et al., , 2017), we avoid using a metric like RMSE that favours (too) narrow posteriors, thus providing misleading interpretations if used as a sole metric.
@@
We summarise our findings using the multi-level model (Bürkner, , 2017) described in Appendix A for both posterior bias and SD. The SE simulation scenario presented in Figure 5 shows that, in most cases, HSGPs have lower bias compared to exact GPs and VIGPs. The VIGPs consistently show the highest bias. Overall, for the SE data scenario, we see on average about 11% (for higher sample sizes) to about 18% (for lower sample sizes) decrease in posterior bias between VIGPs and HSGPs (with $M=22$ basis functions). Results for the posterior SD are more mixed, with VIGPs showing higher SDs for small sample sizes and about the same SDs for larger sample sizes.
@@
In Figure 6 we show the more challenging simulation scenario. Here, the data generating process had a higher amount of variability in length-scales across the output dimensions. We observe a stronger difference in posterior bias between VIGPs and HSGPs (in favour of HSGPs). We hypothesize that HSGPs are more strongly favoured here as they allow for varying hyperparameters across dimensions, while the VIGPs (at least their Pyro implementation) only support constant hyperparameters.
Among the HSGPs, higher number of basis functions ($M=26\text{ }\text{and}\text{ }30$ compared to $M=22$) result in lower posterior bias for both the simulation scenarios shown in Figures 5 and 6. Combining this result with their corresponding model calibration tests, we recommend using $M=30$ to achieve the highest accuracy for the latent variable estimates. For faster but reliable results, HSGPs with $M=22$ can be used in simpler cases like the SE data scenario. Furthermore, in the remaining scenarios (see Figures 18-21 in Appendix E), HSGPs perform consistently better than their competitors. In all the scenarios, each of the fitted models (exact GPs only for $N=20$) show a consistent decrease in posterior bias and SD as the number of output dimensions $D$ and sample size $N$ increases. We report the estimation results for GP hyperparameters in the Appendix B.
@@ 5 Real-World Case Study
We showcase the application of the HSGPs to real-world scRNA sequencing data by re-analysing cell-cycle data from Mahdessian et al., (2021). Our aim is to estimate the underlying cell ordering known as ”pseudotime”. The dataset comprises of spliced single-cell RNA expression profiles along the cell cycle with their own time sequence for cells that are aligned with the experimental time called ”cell hours”. We choose this dataset because it covers a cyclic process going through three distinct biological phases (G1, G2M and S-ph) depicting substantial variation in gene expressions. Additionally, the provided cell hours is a continuous temporal sequence that can be utilised as a prior guide to estimate pseudotime. For the purpose of this case study, we use the data from all the cells along with a selection of 12 influential genes (see Appendix G for the list) based on the recommendations of Mahdessian et al., (2021). In other words, each sample point corresponds to a single cell and each output dimension corresponds to a single gene, with the value being the gene expressions per cell. Thus, the analysed data comprises of $N=960$ observations and $D=12$ output dimensions.
@@
In Figure 8, we show the posterior estimates of the HSGP hyperparameters along with their posterior 95% CIs. A key observation here is the strong differences in the hyperparameter estimates across the genes (output dimensions). This confirms how crucial it is for the model to assume gene-specific hyperparameters while modelling these biological processes. Modelling varying GP hyperparameters, while supported in our HSGPs, is not possible yet for the (pyro implemented) VIGPs, so we do not include their estimates in Figure 8.
@@ 6 Discussion
We generalise Hilbert space approximations for Gaussian processes (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022) to develop a scalable class of latent variable models for high-dimensional multi-output data. Our HSGPs account for both varying hyperparameters across output dimensions and correlations between output dimensions. They reduce the steep computational complexity of $O(N^{3}D+ND^{2})$ for exact multi-output GPs to $O(NMD+ND^{2})$ for a dataset of $N$ samples and $D$ output dimensions, where $M$ is the number of basis functions of the HSGP. Since $M$ is typically much smaller than $N$, using HSGPs reduce computation times drastically. Through extensive simulations, we investigated the statistical properties of latent GPs in terms of model convergence, uncertainty calibration, and estimation accuracy. In all of the above criteria, the HSGPs performed better than both exact GPs and approximate GPs based on inducing points and variational inference (VIGPs)(Titsias, , 2009; Ahmed et al., , 2019). While VIGPs are faster yet, they provide limited modelling capabilities and are outperformed by HSGPs in terms of model calibration and estimation accuracy. Even though HSGPs are comparably slower, they only need a few minutes even for relatively large datasets such as the one investigated in our real-world case study. This stands in sharp contrast to exact GPs, which would have taken at least several hours, if not more, to fit the same data.
@@ 6.1 Limitations and future research
For GPs with Matern class of covariance functions, the length-scale $\rho$ is directly related to the scale of the input space. In latent variable GPs, this implies strong dependencies between the prior on $\rho$, the range of latent $x$, and the prior measurement SD $s$ (see Sections 3 and 4.2). As a result, specifying all three independently may lead to convergence issues and overall bad model behaviour. Future research should investigate these dependencies more closely to provide better guidelines for prior specification in latent variable GPs.
An important generalization of latent variable GPs concern the inclusion of derivative information, which can further increase estimation accuracy (Mukherjee et al., , 2025). However, such methods have so far only been developed for exact GPs, which hamper their real-world applicability. Extending HSGPs to include derivative information will likely provide a scalable solution. However, such an extension requires spectral densities for composite covariance functions, which to our knowledge is yet to be derived.
@@ Acknowledgments
Appendix
Appendix A: Methods related to simulation study
@@
where $\beta,\text{ }b$ (for $\mu_{resp}$) and $\eta,\text{ }u$ (for $\sigma_{resp}$) are coefficients at the population and group levels with $X$ and $Z$ being the corresponding design matrices. The $s_{\tau}(t_{\tau})$ terms denote smooth functions over covariates $t$ fitted via splines. We use these models to summarize the results from our experiments. In the simulation studies, the population level design matrix $X$ contains covariates representing the approximation method and number of output dimensions along with their interaction terms. Based on our experimental setup, the approximation methods include the exact model, HSGPs (with three choices of basis functions) and VIGPs, thus creating a five level factor variable. We use a three level factor variable depicting the $5$, $10$, and $20$ output dimensions. Through $Z$ we account for the group-level dependency structure in the response induced by fitting multiple models to the same simulated dataset. We include a random intercept over datasets as well as corresponding random slopes of the approximation method choices. Lastly, to capture the non-linear relation of the response to the ground truth $t$, we introduce thin-plate regression spline terms $s_{\mu_{\tau}}(t_{\tau})$ and $s_{\sigma_{\tau}}(t_{\tau})$. The spline terms accounts for any non-linear relationship of the response with respect to the true parameter values $t$. We use the log $\gamma$ scores as our response to summarize the model calibration results in Section 4.4. To summarize the results of latent variable estimation in Section 4.5, we subsequently specify posterior bias and SD as the response.
@@
$\gamma=2\quad\min_{j\in\{1,\dots,J+1\}}(\min\{\text{Bin}(R_{j}\mid J,z_{j}),1-\text{Bin}(R_{j}-1\mid J,z_{j})\}),$ (30)
@@
Appendix C: MCMC convergence diagnostics (additional figures)
@@
Appendix D: Model calibration tests (additional figures)
@@
Appendix E: Latent input estimation (additional figures)
@@
Appendix F: Hyperparamter estimation (additional figures)
@@
Appendix G: Additional info for the case study

### v2

@@ 4.5 Latent variable estimation
For all of the simulation scenarios discussed in Section 4.1, we evaluate the estimation capabilities of the different fitted models specified in Section 4.2. We show model performance in terms of absolute bias $|\text{Bias}(x_{post},x_{true})|$ of the latent posterior estimates $x_{post}$ with respect to their true values $x_{true}$ as well as $\text{SD}(x_{post})$ as a measure of posterior sharpness. Since mean-field VI is known to underestimate the posterior SD (Kucukelbir et al., , 2017), we avoid using a metric like RMSE that favors (too) narrow posteriors, thus providing misleading interpretations if used as a sole metric.
@@
We summarize our findings using the multi-level model (Bürkner, , 2017) described in supplementary materials Section A for both posterior bias and SD. The SE simulation scenario presented in Figure 5 shows that, in most cases, HSGPs have lower bias compared to exact GPs and VIGPs. The VIGPs consistently show the highest bias. Overall, for the SE data scenario, we see on average about 11% (for higher sample sizes) to about 18% (for lower sample sizes) decrease in posterior bias between VIGPs and HSGPs (with $M=22$ basis functions). Results for the posterior SD are more mixed, with VIGPs showing higher SDs for small sample sizes and about the same SDs for larger sample sizes.
@@
In Figure 6 we show the more challenging simulation scenario. Here, the data generating process had a higher amount of variability in length-scales across the output dimensions. We observe a stronger difference in posterior bias between VIGPs and HSGPs (in favor of HSGPs). We hypothesize that HSGPs are more strongly favored here as they allow for varying hyperparameters across dimensions, while the VIGPs (at least their Pyro implementation) only support constant hyperparameters.
Among the HSGPs, higher number of basis functions ($M=26\text{ }\text{and}\text{ }30$ compared to $M=22$) result in lower posterior bias for both the simulation scenarios shown in Figures 5 and 6. Combining this result with their corresponding model calibration tests, we recommend using $M=30$ to achieve the highest accuracy for the latent variable estimates. For faster but reliable results, HSGPs with $M=22$ can be used in simpler cases like the SE data scenario. Furthermore, in the remaining scenarios (see Figures S10-S13 in supplementary materials Section E), HSGPs perform consistently better than their competitors. In all the scenarios, each of the fitted models (exact GPs only for $N=20$) show a consistent decrease in posterior bias and SD as the number of output dimensions $D$ and sample size $N$ increases. We report the estimation results for GP hyperparameters in the supplementary materials Section B.
@@ 5 Real-World Case Study
We showcase the application of the HSGPs to real-world scRNA sequencing data by re-analyzing cell-cycle data from Mahdessian et al., (2021). Our aim is to estimate the underlying cell ordering known as ”pseudotime”. The dataset comprises of spliced single-cell RNA expression profiles along the cell cycle with their own time sequence for cells that are aligned with the experimental time called ”cell hours”. We choose this dataset because it covers a cyclic process going through three distinct biological phases (G1, G2M and S-ph) depicting substantial variation in gene expressions. Additionally, the provided cell hours is a continuous temporal sequence that can be utilized as a prior guide to estimate pseudotime. For the purpose of this case study, we use the data from all the cells along with a selection of 12 influential genes (see supplementary materials Section G for the list) based on the recommendations of Mahdessian et al., (2021). In other words, each sample point corresponds to a single cell and each output dimension corresponds to a single gene, with the value being the gene expressions per cell. Thus, the analyzed data comprises of $N=960$ observations and $D=12$ output dimensions.
@@
In Figure 8, we show the posterior estimates of the HSGP hyperparameters along with their posterior 95% CIs. A key observation here is the strong differences in the hyperparameter estimates across the genes (output dimensions). This confirms how crucial it is for the model to assume gene-specific hyperparameters while modeling these biological processes. Modeling varying GP hyperparameters, while supported in our HSGPs, is not possible yet for the (pyro implemented) VIGPs, so we do not include their estimates in Figure 8.
@@ 6 Discussion
We generalize Hilbert space approximations for Gaussian processes (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022) to develop a scalable class of latent variable models for high-dimensional multi-output data. Our HSGPs account for both varying hyperparameters across output dimensions and correlations between output dimensions. They reduce the steep computational complexity of $O(N^{3}D+ND^{2})$ for exact multi-output GPs to $O(NMD+ND^{2})$ for a dataset of $N$ samples and $D$ output dimensions, where $M$ is the number of basis functions of the HSGP. Since $M$ is typically much smaller than $N$, using HSGPs reduce computation times drastically. Through extensive simulations, we investigated the statistical properties of latent GPs in terms of model convergence, uncertainty calibration, and estimation accuracy. In all of the above criteria, the HSGPs performed better than both exact GPs and approximate GPs based on inducing points and variational inference (VIGPs)(Titsias, , 2009; Ahmed et al., , 2019). While VIGPs are faster yet, they provide limited modeling capabilities and are outperformed by HSGPs in terms of model calibration and estimation accuracy. Even though HSGPs are comparably slower, they only need a few minutes even for relatively large datasets such as the one investigated in our real-world case study. This stands in sharp contrast to exact GPs, which would have taken at least several hours, if not more, to fit the same data.
@@ 6.1 Limitations and future research
For GPs with Matern class of covariance functions, the length-scale $\rho$ is directly related to the scale of the input space. In latent variable GPs, this implies strong dependencies between the prior on $\rho$, the range of latent $x$, and the prior measurement SD $s$ (see Sections 3 and 4.2). As a result, specifying all three independently may lead to convergence issues and overall bad model behavior. Future research should investigate these dependencies more closely to provide better guidelines for prior specification in latent variable GPs.
With regard to the correlation structure $C$ across output dimensions (see Section 3.3.1), we make the choice of modeling it on the level of the GP functions $f$ rather than on the level of observations $y$. It is not always apparent which of these two options is preferable for a given application. For example, in spatial statistics, correlations between discrete locations are also either modeled on the mean (similar to the GP functions in our case) level, e.g., via conditional autoregressive (CAR) structures, or on the observation level, e.g., via spatial autoregressive (SAR) structures (Wall, , 2004). In this paper, we apply our methods to RNA gene expressions for cell cycle data, where the genes are known to be correlated at the biological process level (Riba et al., , 2022). Thus, specifying across-dimension correlations on the GP functions appear sensible here. For other applications, it may be preferable to specify the correlations at the observational level, or even jointly at both levels.
Another consideration concerning $C$ is whether to model it as input dependent or input independent. In this paper, we chose it to be input independent, thus estimating the full correlation matrix directly. As an alternative, an input dependent formulation could be achieved through another GP that models the correlations across output dimensions within each observation (similar to a cross-covariance structure shown in Gneiting et al., (2010)). In this regard, it would further be interesting to study Hilbert space approximations also for this across-output dimension GP, especially when the number of dimensions is large enough for exact GPs to become computationally impractical. We leave these extensions of our framework to future research.
Lastly, an important generalization of latent variable GPs concern the inclusion of derivative information, which can further increase estimation accuracy (Mukherjee et al., , 2025). However, such methods have so far only been developed for exact GPs, which hamper their real-world applicability. Extending HSGPs to include derivative information will likely provide a scalable solution. However, such an extension requires spectral densities for composite covariance functions, which to our knowledge is yet to be derived.
@@ Acknowledgments
Supplementary materials
A: Methods related to simulation study
@@
where $\beta,\text{ }b$ (for $\mu_{resp}$) and $\eta,\text{ }u$ (for $\sigma_{resp}$) are coefficients at the population and group levels with $X$ and $Z$ being the corresponding design matrices. The $s_{\tau}(t_{\tau})$ terms denote smooth functions over covariates $t$ fitted via splines. We use these models to summarize the results from our experiments. In the simulation studies, the population level design matrix $X$ contains covariates representing the approximation method and number of output dimensions along with their interaction terms. Based on our experimental setup, the approximation methods include the exact model, HSGPs (with three choices of basis functions) and VIGPs, thus creating a five level factor variable. We use a three level factor variable depicting the $5$, 1010, and $20$ output dimensions. Through $Z$ we account for the group-level dependency structure in the response induced by fitting multiple models to the same simulated dataset. We include a random intercept over datasets as well as corresponding random slopes of the approximation method choices. Lastly, to capture the non-linear relation of the response to the ground truth $t$, we introduce thin-plate regression spline terms $s_{\mu_{\tau}}(t_{\tau})$ and $s_{\sigma_{\tau}}(t_{\tau})$. The spline terms accounts for any non-linear relationship of the response with respect to the true parameter values $t$. We use the log $\gamma$ scores as our response to summarize the model calibration results in Section 4.4. To summarize the results of latent variable estimation in Section 4.5, we subsequently specify posterior bias and SD as the response.
@@
$\gamma=2\qquad\min_{j\in\{1,\dots,J+1\}}(\min\{\text{Bin}(R_{j}\mid J,z_{j}),1-\text{Bin}(R_{j}-1\mid J,z_{j})\}),$ (30)
@@
C: MCMC convergence diagnostics (additional figures)
@@
D: Model calibration tests (additional figures)
@@
E: Latent input estimation (additional figures)
@@
F: Hyperparamter estimation (additional figures)
@@
G: List of genes for the case study